Please answer the following question:
The sum of their ages is 20. How old is RiRicardo is 4 years less than twice as old as his sister.cardo? Thanks a million!

Answer by smoothflyin(6)

To answer this equation you really just need to assign variables to the various elements. So for instance, we could say that
R=Ricardo's age
S=His sister's age
Ricardo is 4 years less than twice as old as his sister
So
R=2S-4
and we know that the sum of their ages is equal to 20, so
(R+S=20) and using the commutative property
(S=20-R)
So now we plug it back into our original problem (R=2S-4) and we get
R=2(20-R)-4
R=40-2R-4
we bring to 2R over to the other side of the equal sign and change its sign
3R=40-4 or 3R=36
divide through by 3 and you get R=12
Ricardo's age is 12 and his sister's age is 8...Hope this helped!